OpenStudy (anonymous):

X (SQUARED)5 DIVIDED BY X (SQUARED )-8

hartnn (hartnn):

is it, \[\frac{x^5}{x^{-8}}\] or \[\frac{x^2*5}{x^2-8} \] ??

OpenStudy (anonymous):

the top one

OpenStudy (zzr0ck3r):

x^(5-(-8)) = x^13

hartnn (hartnn):

use this: \[\frac{x^a}{x^b}=x^{a-b}\] does this help ?

OpenStudy (anonymous):

ok thank you how would you simplify xy (squared)-6

OpenStudy (anonymous):

\[\frac{x^a}{x^{-b}} \implies x^{a +b}\]

OpenStudy (anonymous):

yes that helps

OpenStudy (anonymous):

\[\frac{x^a}{x^{-b}}=x^{a+b} \]

OpenStudy (zzr0ck3r):

First only say squared when there is a 2 up there. use the carrot ^ so x squared is x^2 and x to the fifth power is x^5

OpenStudy (zzr0ck3r):

so your asking about (x*y)^(-6)?

OpenStudy (zzr0ck3r):

= x^(-6)*y^(-6)

hartnn (hartnn):

\[(xy)^{-6}\] right ? \[(xy)^a=x^a*y^a\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(xy) -6

OpenStudy (anonymous):

xy^-6

OpenStudy (zzr0ck3r):

make sure you use the brackets accordingly (xy)^(-6) = x^(-6) * y^(-6) but does not equal xy^(-6) ( this is the same thing as x * y^(-6))